C=450+0.4Y
I=350
G=150
X=70
Z=35+0,1Y
T=0.15Y
Yf=1550

Calculate the level of autonomous spending in this economy
1
Expert's answer
2020-12-01T10:19:49-0500

Consumption: C = 450 + 0.4Y where Y is income

Investment: I = 350

Government spending: G = 150

Exports: X = 70

Imports: Z = 35 + 0.1Y

Taxes: T = 0.15Y

Full employment income: Yf = 1550

Y = C + I + G + X – Z

Y = 450 + 0.4Y + 350 + 150 + 70 – 35 – 0.1Y

Y = 985 + 0.4(Y – T) – 0.1Y

Y = 985 + 0.4(Y – 0.15Y) – 0.1Y

Y = 985 + 0.24Y

Autonomous Spending = 985